IODP Proceedings    Volume contents     Search



The primary objective for Hole U1382A was to install a single-level CORK to perform long-term coupled microbiological, biogeochemical, and hydrological experiments in uppermost basaltic crust. We installed a reentry cone with 53 m of 16 inch casing, installed 102 m of 10¾ inch casing and cemented it a few meters into the basaltic basement, cored 100 m of basalt below the casing, conducted downhole logging and hydrologic (packer) experiments, and installed a CORK that extends to 188.7 mbsf.

After operations in Hole 395A, we intended to go directly to Hole U1382A (50 m west of Hole 395A). However, we had to leave the area because of Tropical Storm Philippe. After the ship was secured for transit at 0215 h on 29 September 2011, we headed to the northeast to avoid the approaching storm (all times are local ship time, Universal Time Coordinated [UTC] – 3 h). After Tropical Storm Philippe crossed over the North Pond drilling area, we returned to Hole U1382A at 1224 h on 1 October. All operational tasks at Site U1382, along with task start and end times, are listed in Table T1.

We assembled a reentry cone and 52.98 m of 16 inch casing and lowered it to the seafloor. The trip was temporarily suspended to install the camera system, which was lowered as the pipe trip continued to just above the seafloor. When the casing shoe was just above the seafloor, the top drive was picked up and the drill string was spaced out to start Hole U1382A. We started jetting in the reentry system at 0745 h on 2 October. After 2.75 h the reentry cone mud skirt was landed on the seafloor and verified with the camera system, and the casing running tool was released from the casing. The camera system was then pulled to the surface as the drill string and running tool were pulled back to the rig floor. The bottom-hole assembly (BHA) was set back, and the running tool was detorqued at the rig floor. The BHA for drilling the 14¾ inch hole was assembled, and the drill string was tripped to just above the reentry system. During the pipe trip, the camera system was installed and lowered as the drill string was tripped to bottom. After the drill string was spaced out for reentry, the bit reentered Hole U1382A at 0348 h on 3 October. The top drive was then picked up, and the drill string was run to the casing shoe and spaced out for drilling. The sediment section was drilled without coring from 4547 to 4584 meters below rig floor (53–90 mbsf), at which point basement was contacted. Drilling parameters indicated a fairly hard formation from 90 to 93 mbsf, but drilling proceeded fairly quickly from 93 to 99 mbsf. From 99 to 110 mbsf, drilling parameters again were slow and consistent, indicating a hard formation. We decided to terminate the hole at 110 mbsf to allow 8 m of rathole below 10¾ inch casing that would extend to 102 mbsf. After the hole was conditioned, the bit was tripped above the casing shoe, the top drive was removed, and the drill string was tripped back to the rig floor. Before we could begin rigging up for running casing, we had to slip and cut 115 ft of drill line. We assembled 101.86 m of 10¾ inch casing that included a 14¾ inch outside diameter swellable packer one joint from the top and a casing hanger with a seal ring. We attached the casing running tool, lowered the entire casing string to the seafloor, and reentered Hole U1382A at 1355 h on 4 October. The casing was lowered smoothly into the hole until the last couple of meters, when we had to circulate to help clear the way so it could fully land. After the casing string was fully landed and latched, we cemented it in place with 20 bbl of cement blended with lost-circulation material (Cello-Flake). The cement displacement calculation was made to leave ~15 m of cement inside the casing above the casing shoe. Once we released the casing running tool from the casing hanger (1712 h on 4 October), we pumped seawater through the drill string to clean it of any remaining cement. Before we retrieved the casing running tool, we performed a 30 min survey of the Hole 395A reentry system (50 m to the west). The casing running tool arrived back on the rig floor at 0200 h on 5 October.

Our next step was RCB coring, so we assembled a new C-7 RCB bit (with center bit) with a three-stand BHA and lowered it to the seafloor. A break in tripping pipe occurred around 1030 h to install the camera system, but the system was quickly retrieved when the subsea camera did not work. The pipe trip continued for another hour and was halted again to install the repaired camera system. At 1330 h the drill string was spaced out for reentry, and Hole U1382A was quickly reentered at 1337 h. The bit was carefully run into the hole, and cement was encountered 14 m above the casing shoe (1 m below the expected 15 m). The top drive was then picked up, and the cement was drilled from 88 mbsf to just below the casing shoe that had been positioned at 102 mbsf. We circulated mud to clean the hole, recovered the center bit by wireline, dropped an RCB core barrel, and RCB coring began. The first core on deck, Core 336-U1382A-2R, arrived at 2245 h on 5 October. Coring continued through Core 336-U1382A-12R, which was on board at 2320 h on 7 October. We cored 100 m of section from 110 to 210 mbsf and recovered 31.8 m (32%; Table T2). After coring was completed, five wiper trips were made from total depth to the casing shoe and back to total depth. The first three trips revealed tight spots and circulation problems. The fourth and fifth trips were made with no evidence of drag or circulation problems, and no fill was found at the bottom of the hole. After hole cleaning and conditioning were completed, the drill string was tripped from the hole, clearing the seafloor at 0729 h on 8 October.

After the bit was back on the rig floor and before logging began, we assembled a stand of 6¾ inch perforated and coated drill collars for the lowermost portion of the CORK installation. We wanted to make up the drill collars before assembling the CORK to make it easier and more efficient to paint with epoxy prior to being deployed. Next, we assembled a logging BHA with a logging bit and the drill string packer, lowered it 64 m into Hole U1382A, and began deploying the logging tools. Logging proceeded with a modified triple combination tool string, with the DEBI-t on the bottom of the string (see “Downhole logging”). Log data were collected while the string was lowered to the bottom of the hole. However, while logging upward, the power failed ~20 m below the casing. The tool string was pulled to the surface, and the problem was found to be in the cablehead. The cablehead was reterminated, and we decided to run the FMS-sonic tool string next. After two successful passes, the lower portion of the FMS (calipers) would not enter the logging bit. After 2 h of working the string up and down, opening and closing the calipers, and pumping seawater, the entire tool string was able to pass through the logging bit; when it was recovered, one of the caliper arms had been damaged. We then spaced out the drill string for the hydrologic (packer) flow test and began attempting to inflate the packer. We made four attempts to set the packer inside the 10¾ inch casing; however, each time, high vessel heave (3 m) caused the packer to deflate, so the experiment was terminated. Before pulling out of the hole and installing the CORK, we lowered the entire BHA (including the deflated packer) until the logging bit reached the bottom of the hole to check that the hole was still open to full depth. We did not encounter any problem intervals and found only ~1.5 m of fill. The drill string was recovered, and the bit was back on board at 1328 h on 10 October. Before we could begin our next operation, we had to slip and cut 115 ft of drill line. We started assembling the CORK at 1500 h on 10 October. Details of the CORK are shown in Figures F1 and F2 and Tables T3 and T4.

The preassembled 6¾ inch coated, painted, and perforated drill collar stand was picked up and run through the rotary table. Epoxy paint was used to touch up rust marks, and 10% ethanol was used to wipe grease from all exposed steel between the bottom of the casing and the top of the combination packer. After a crossover was installed, we attached 15.35 m of 5½ inch coated and perforated casing. Miniscreens were attached to the outside of the lowermost perforated 5½ inch casing joint. The lowermost four miniscreens included one titanium microbiology screen (connected to a Tefzel umbilical) and three stainless steel miniscreens for chemistry (two connected to ⅛ inch stainless umbilicals and one attached to a ¼ inch umbilical). The next set of four miniscreens, including a second titanium microbiology screen (connected to a Tefzel umbilical) and three stainless steel miniscreens for chemistry (one connected to a ⅛ inch stainless umbilical and two attached to ¼ inch umbilicals), was placed just above the first set of miniscreens. The stainless steel miniscreen for pressure (connected to a ¼ inch stainless umbilical) was strapped to the casing immediately above the second set of miniscreens. Crossovers to the 4½ inch fiberglass casing were installed, and then 44.25 m of 4½ inch fiberglass casing was made up. Umbilicals from the miniscreens and plastic Kwik-zip centralizers were installed on the outside of the casing as it was run. Next to be installed was a crossover from the fiberglass casing to the landing seat for the instrument string, followed by the joint with the swellable and inflatable packers. Umbilicals were terminated and connected to the bottom and top of the combination packer. Above the packers, ~91.01 m of uncoated and unperforated 4½ inch steel casing was run, followed by the CORK head. All pressure and sampling lines were connected to the bottom of the CORK. It took us 10 h to complete assembly, and the bottom of the CORK string extends to 188.7 mbsf.

Our next step was to assemble and install the OsmoSampler string into the CORK at the rig floor (see “CORK observatory”). Modifications were made before and during installation (thicker springs were installed, and parts were ground off of the latching mechanisms so they would stick out further) so that it would latch into the CORK head; however, these modifications did not work. The CORK head was submerged for 10 min to clear the pressure line of air, and then the CORK head was raised through the moonpool. A fast-flow OsmoSampler system was attached, and all of the unused valves were closed. The used valves include one pressure valve and a ⅛ inch chemistry line attached to the fast-flow OsmoSampler (see “CORK observatory”). The CORK was then lowered to ~100 m. The camera system was test fit over the CORK, and at 0630 h on 11 October, we began lowering the CORK package to the seafloor. The camera system was installed, and Hole U1382A was reentered for the last time at 1630 h on 11 October.

After carefully lowering the CORK into the hole, we landed the CORK at 1820 h on 11 October. The packer was inflated (to 1400 psi, which then bled off to ~1280–1300 psi), and the camera was pulled to the surface. Next, the remotely operated vehicle (ROV) platform was assembled and rigged up with the ROV deployment tool attached below the camera system and then lowered back to the seafloor. The platform was released over the CORK at 0055 h on 12 October. The camera system was retrieved and after the ROV deployment system was removed, the camera was lowered back to the bottom to observe the final step of installing the Hole U1382A CORK—releasing the running tool from the CORK. This was successfully released at 0425 h on 12 October. The drill string with the CORK running tool was recovered back on board at 1145 h, ending Hole U1382A.